OpenStudy (anonymous):

An airplane flew 3 hours with a 60 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air.

OpenStudy (dumbcow):

d=rt set distances equal 3(r-60) = 2(r+60) 3r -180 = 2r +120 r = 300